OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)?

OpenStudy (anonymous):

Distance formula = sqrt ( (y2 - y1)^2 + (x2-x2)^2)

OpenStudy (anonymous):

d = sqrt( (3--2)^2 + (7-1)^2 ) = sqrt(37)

OpenStudy (anonymous):

= 6.08

OpenStudy (anonymous):